# Mixing Time Bounds for the Gibbs Sampler under Isoperimetry

### v1

@@ Abstract
We establish bounds on the conductance for the systematic-scan and random-scan Gibbs samplers when the target distribution satisfies a Poincaré or log-Sobolev inequality and possesses sufficiently regular conditional distributions. These bounds lead to mixing time guarantees that extend beyond the log-concave setting, offering new insights into the convergence behavior of Gibbs sampling in broader regimes. Moreover, we demonstrate that our results remain valid for log-Lipschitz and log-smooth target distributions. Our approach relies on novel three-set isoperimetric inequalities and a sequential coupling argument for the Gibbs sampler.
@@ 1 Introduction
Sampling from a target distribution has long been a central problem in many fields related to Statistics and Applied Probability. Markov chain Monte Carlo (MCMC) methods are among the most popular algorithms for generating samples from a potentially high-dimensional target distribution $\pi$. The idea is to simulate an ergodic Markov chain that admits $\pi$ as its stationary distribution, and use the states of the chain as approximate samples from $\pi$. The Gibbs sampler (Geman and Geman, 1984) is one of the most widely used MCMC methods in practical applications. This algorithm updates a single coordinate (or a block of coordinates) at each iteration using its conditional distribution under $\pi$, while keeping the other coordinates fixed. There are two common variants: systematic-scan Gibbs and random-scan Gibbs. In the systematic-scan case, coordinates are updated one by one in a fixed order. In random-scan Gibbs, one coordinate is chosen randomly at each step to be updated. The advantage of Gibbs sampling is that it simplifies high-dimensional sampling by reducing it to a series of lower-dimensional updates. Moreover, in many popular settings, such as hierarchical or exponential family models, the conditional distributions are tractable, eliminating the need for step-size selection.
@@
In this paper, we characterize the convergence of Gibbs sampling using the so-called mixing time, which quantifies the number of iterations needed for the Markov chain to be within a specified distance of its stationary distribution. A common technique for bounding the mixing time is to analyze a quantity known as the conductance of the Markov chain. One of the key components of this technique relies on the notion of isoperimetry for the target distribution. This is often formalized via Cheeger’s isoperimetric inequality (Cheeger, 1970), which holds for log-concave target distributions (Bobkov, 1999) and remains valid under bounded perturbations, as characterized by the criterion of Holley and Stroock (1987). The conductance approach has proven successful for the Hit-and-Run algorithm (Lovász and Simonovits, 1993; Lovász, 1999; Lovász and Vempala, 2004; Vempala, 2005), as well as for random-walk Metropolis-type algorithms (Dwivedi et al., 2019; Mangoubi and Vishnoi, 2019; Chen et al., 2020; Chewi et al., 2021; Zou et al., 2021; Wu et al., 2022; Mou et al., 2022; Chen and Gatmiry, 2023), among others. However, conductance-based proof techniques have seen limited use in the context of Gibbs sampling, with notable exceptions including sampling from uniform distributions over convex bodies (Narayanan and Srivastava, 2022; Laddha and Vempala, 2023), posterior distributions in Bayesian regression models (Lee and Zhang, 2024), and strongly log-concave targets (Wadia, 2024).
@@
In the recent seminal work by Andrieu et al. (2024), the authors unify many ideas from the references mentioned above and provide clear guidelines for controlling the mixing time of general MCMC methods under the assumption of isoperimetry for the target distribution. Specifically, they consider log-concave target distributions which also satisfy either a Poincaré or a log-Sobolev inequality. However, log-concavity is not necessary for these functional inequalities to hold: a wide range of non-log-concave distributions still satisfy a Poincaré or log-Sobolev inequality, making the class of distributions satisfying such functional inequalities substantially broader. Nevertheless, when deriving explicit bounds on the mixing time of the random-walk Metropolis algorithm, even these authors exclusively focus on the strongly log-concave case. To the best of our knowledge, the work of Vempala and Wibisono (2019) on the unadjusted Langevin algorithm is the only existing result that establishes convergence bounds under a Poincaré or log-Sobolev inequality rather than relying on Cheeger’s isoperimetric inequality. It is therefore interesting to investigate whether the framework developed by Andrieu et al. (2024) can be applied to the Gibbs sampler targeting non-log-concave distributions, specifically those satisfying either a Poincaré or log-Sobolev inequality.
@@ 1.1 Our contributions
In this paper, we establish conductance bounds for the systematic-scan and random-scan Gibbs samplers over a broad class of target distributions supported on $\mathbb{R}^{d}$. In particular, consider a target distribution $\pi$ satisfying a $L^{q}$-Poincaré inequality with constant $\mathtt{p}_{q}$ and whose one-dimensional conditional distributions are Hölder continuous in total variation distance with constant $M$ and exponent $\beta$ (see Subsection 2.2 for formal definitions). Denote by $P_{\mathrm{SS}}$ and $P_{\mathrm{RS}}$ the Markov kernels corresponding to the systematic-scan and random-scan Gibbs samplers, respectively. Then, for sufficiently large $d$, our main results (Theorem 4.7 and Corollary 4.9) provide the following bounds on the conductance of the systematic-scan Gibbs kernel, $\Phi(P_{\mathrm{SS}})$, and the random-scan Gibbs kernel, $\Phi(P_{\mathrm{RS}})$:
$\Phi(P_{\mathrm{SS}})\gtrsim\frac{\mathtt{p}_{q}}{\left(Md\right)^{q/\beta}},$
$\Phi(P_{\mathrm{RS}})\gtrsim\frac{\mathtt{p}_{q}}{M^{q/\beta}(d\cdot\log d)^{q/\beta+1}},$
where $\gtrsim$ denotes inequalities that hold up to a universal constant. Bounds on the mixing time in total variation distance then follow directly from Lovász and Simonovits (1993, Corollary 1.5).
@@
  • We prove new three-set isoperimetric inequalities for distributions satisfying a $L^{q}$-Poincaré (or log-Sobolev) inequality, without assuming log-concavity (Propositions 4.1 and 4.3). This is achieved by applying these functional inequalities to a carefully chosen test function.
@@
  • We show that if $\pi$ is log-Lipschitz continuous or log-smooth, then the one-dimensional conditional distributions under $\pi$ are Hölder continuous in total variation distance (Lemmas 5.4 and 5.5). The proof relies on the fact that log-Lipschitz continuity and log-smoothness are inherited by the marginal distributions of $\pi$, as shown in Lemma 5.3.
@@ 1.2 Connections with recent work
As a special case, the conductance approach has been successfully applied to the random-scan Gibbs sampler targeting uniform distributions over convex bodies, where quantitative convergence guarantees have been established (Narayanan and Srivastava, 2022; Laddha and Vempala, 2023). Narayanan and Srivastava (2022) establish a close coupling condition for a multi-step Metropolis-within-Gibbs scheme, where the corresponding Markov kernel updates a randomly chosen coordinate using a single step of random-walk Metropolis while keeping the others fixed. They then compare the conductance of this auxiliary Markov chain with that of the original random-scan Gibbs sampler. Laddha and Vempala (2023) take a different approach, proving a three-set isoperimetric inequality for partitions of space into axis-disjoint regions. This allows them to control the mixing behavior without relying on auxiliary chains. In our work, we extend these ideas to general target distributions and establish three-set isoperimetric inequalities that hold for arbitrary partitions.
Additionally, the conductance approach has been used to establish rapid mixing of the Gibbs sampler targeting the posterior distributions in Bayesian probit, logit, and lasso regression models (Lee and Zhang, 2024). In each case, the Gibbs sampler is made tractable by augmenting the state space with latent variables. For such data-augmented chains, Lee and Zhang (2024) show that close coupling conditions follow directly from the conditional independence structure of the posterior distribution. Isoperimetric inequalities are obtained via transference inequalities to a transformed Markov chain with a log-concave target. Under suitable assumptions on the data-generating process, these authors then establish bounds on the mixing time that hold with high probability.
@@ 1.3 Organization
The rest of this paper is structured as follows. Section 2 introduces the Gibbs sampler and outlines the relevant assumptions on the target distribution. Section 3 reviews standard definitions and key results on Markov chain mixing times. In particular, we explain how a three-set isoperimetric inequality and a close coupling condition can be used to derive lower bounds on conductance. Next, we establish three-set isoperimetric inequalities for target distributions satisfying a Poincaré or log-Sobolev inequality (Subsection 4.1) and derive close coupling conditions for the systematic-scan and random-scan Gibbs kernels (Subsection 4.2). We then use these results to prove conductance and mixing time bounds for both variants of the Gibbs sampler (Subsection 4.3). Finally, Section 5 extends these bounds to log-Lipschitz continuous and log-smooth target distributions. We also provide examples of non-log-concave target distributions for which our results can be applied to establish mixing time bounds for both variants of the Gibbs sampler, filling a gap not addressed by any other existing literature.
@@ 2.1 The Gibbs sampler
We consider a probability distribution $\pi(\mathrm{d}x)$ supported on $\mathbb{R}^{d}$, equipped with its standard Borel $\sigma$-algebra. Let $P(x,\mathrm{d}y)$ be the Markov transition kernel of a discrete-time Markov chain on $\mathbb{R}^{d}$ with stationary (or invariant) distribution $\pi$. Here, the Markov chain is interpreted as the MCMC method for generating approximate samples from $\pi$. The kernel $P$ is said to be reversible with respect to $\pi$ if
@@
Reversibility with respect to $\pi$ implies stationarity of $\pi$, however the reverse implication is in general false. Among MCMC methods, we focus on the Gibbs sampler. This method has two commonly used variants: the random-scan Gibbs sampler and the systematic-scan Gibbs sampler, which we now introduce. Let $x=\left(x_{1},\dots,x_{d}\right)\in\mathbb{R}^{d}$. For $1\leq k\leq l\leq d$, we write
@@
We additionally use the notation $\left[d\right]=\left\{1,\dots,d\right\}$. Let $X\sim\pi$. Denote the conditional distribution of $X_{k}$ given $X_{-k}=x_{-k}$ as $\pi(\cdot\mid x_{-k})$. For all $k\in[d]$, define the Markov kernel
$P_{k}(x,\mathrm{d}y)=\pi(\mathrm{d}y_{k}\mid x_{-k})\delta_{x_{-k}}(\mathrm{d}y_{-k}),$
where $\delta_{x_{-k}}$ denotes the $(d-1)$-dimensional Dirac delta distribution: for all measurable $S\subseteq\mathbb{R}^{d-1}$,
$\delta_{x_{-k}}(S)=\begin{cases}1&\text{if }x_{-k}\in S,\\ 0&\text{if }x_{-k}\notin S.\end{cases}$
Using the kernels $P_{k}$, we now define the Markov kernels corresponding to the random-scan and systematic-scan Gibbs samplers. For two Markov kernels $P$ and $Q$, we denote their composition as $PQ(x,\mathrm{d}y)=\int P(u,\mathrm{d}y)Q(x,\mathrm{d}u)$.
@@
The Markov kernel associated with the random-scan Gibbs sampler is defined as the mixture of the Markov kernels $P_{k}$:
@@
The Markov kernel associated with the systematic-scan Gibbs sampler is defined as the composition of the Markov kernels $P_{k}$:
$P_{\mathrm{SS}}=P_{d}P_{d-1}\dots P_{2}P_{1}.$
@@
It is worth noting that the Markov kernel associated with the random-scan Gibbs sampler is $\pi$-reversible, whereas the Markov kernel associated with the systematic-scan Gibbs sampler is generally not. To address this, one approach is to use the multiplicative reversibilization of the systematic-scan Gibbs sampler. The kernel for this version is defined as
$P_{\mathrm{SS}}^{*}P_{\mathrm{SS}}=P_{1}P_{2}\dots P_{d-1}P_{d}P_{d-1}\dots P_{2}P_{1},$
where $P^{*}$ represents the adjoint operator of a Markov kernel $P$. In this version, the coordinates are updated first in increasing order, from $1$ to $d$, and then in decreasing order, from $d$ to $1$.

### v2

@@ Abstract
We establish bounds on the conductance for the systematic-scan and random-scan Gibbs samplers when the target distribution satisfies a Poincaré or log-Sobolev inequality and possesses sufficiently regular conditional distributions. These bounds lead to mixing time guarantees that extend beyond the log-concave setting, offering new insights into the convergence behavior of Gibbs sampling in broader regimes. Moreover, we demonstrate that our results remain valid for log-Lipschitz and log-smooth target distributions. Our approach relies on novel isoperimetric inequalities and a sequential coupling argument for the Gibbs sampler.
@@ 1 Introduction
Sampling from a target distribution has long been a central problem in many fields related to statistics and applied probability. Markov chain Monte Carlo (MCMC) methods are among the most popular algorithms for generating samples from a potentially high-dimensional target distribution $\pi$. The idea is to simulate an ergodic Markov chain that admits $\pi$ as its invariant distribution, and use the states of the chain as approximate samples from $\pi$. The Gibbs sampler (Geman and Geman, 1984) is one of the most widely used MCMC methods in practical applications. This algorithm updates a single coordinate (or a block of coordinates) at each iteration using its conditional distribution under $\pi$, while keeping the other coordinates fixed. There are two common variants: systematic-scan Gibbs and random-scan Gibbs. In the systematic-scan case, coordinates are updated one by one in a fixed order. In random-scan Gibbs, one coordinate is chosen randomly at each step to be updated. The advantage of Gibbs sampling is that it simplifies high-dimensional sampling by reducing it to a series of lower-dimensional updates. Moreover, in many popular settings, such as hierarchical or exponential family models, the conditional distributions are tractable, eliminating the need for step-size selection.
@@
In this paper, we characterize the convergence of Gibbs sampling using the so-called mixing time, which quantifies the number of iterations needed for the Markov chain to be within a specified distance of its invariant distribution. A common technique for bounding the mixing time is to analyze a quantity known as the conductance of the Markov chain. One of the key components of this technique relies on the notion of isoperimetry for the target distribution. This is often formalized via Cheeger’s isoperimetric inequality (Cheeger, 1970), which holds for log-concave target distributions (Bobkov, 1999) and remains valid under bounded perturbations, as characterized by the criterion of Holley and Stroock (1987). The conductance approach has proven successful for the Hit-and-Run algorithm (Lovász and Simonovits, 1993; Lovász, 1999; Lovász and Vempala, 2004; Vempala, 2005), as well as for random-walk Metropolis-type algorithms (Dwivedi et al., 2019; Mangoubi and Vishnoi, 2019; Chen et al., 2020; Chewi et al., 2021; Zou et al., 2021; Wu et al., 2022; Mou et al., 2022a; Chen and Gatmiry, 2023), among others. However, conductance-based proof techniques have seen limited use in the context of Gibbs sampling, with notable exceptions including sampling from uniform distributions over convex bodies (Narayanan and Srivastava, 2022; Laddha and Vempala, 2023; Narayanan et al., 2024), posterior distributions in Bayesian regression models (Lee and Zhang, 2024), and strongly log-concave targets (Wadia, 2024).
@@
In the recent seminal work by Andrieu et al. (2024), the authors unify many ideas from the references mentioned above and provide clear guidelines for controlling the mixing time of general MCMC methods under the assumption of isoperimetry for the target distribution. Specifically, they consider log-concave target distributions which also satisfy either a Poincaré or a log-Sobolev inequality. However, log-concavity is not necessary for these functional inequalities to hold: a wide range of non-log-concave distributions still satisfy a Poincaré or log-Sobolev inequality, making the class of distributions satisfying such functional inequalities substantially broader. Nevertheless, when deriving explicit bounds on the mixing time of the random-walk Metropolis algorithm, even these authors exclusively focus on the strongly log-concave case.
For several other classes of sampling algorithms, convergence bounds under a Poincaré or log-Sobolev inequality have been established without relying on Cheeger’s isoperimetric inequality. These include discretizations of the overdamped Langevin diffusion (Altschuler and Chewi, 2024b; Chewi et al., 2025; Mou et al., 2022b; Balasubramanian et al., 2022; Kandasamy and Nagaraj, 2024; Mitra and Wibisono, 2025; Vempala and Wibisono, 2019; Anari et al., 2024), discretizations of the underdamped Langevin diffusion (Altschuler et al., 2025; Lehec, 2025; Ma et al., 2021; Zhang et al., 2023; Camrud et al., 2023; Kandasamy and Nagaraj, 2024; Anari et al., 2024), as well as proximal samplers (Wibisono, 2019; Chen et al., 2022; Fan et al., 2023; Liang and Chen, 2022; Mitra and Wibisono, 2025), among others. In these works, convergence is typically established by comparing the discrete-time sampling algorithm to an associated continuous-time process, such as the Langevin diffusion, whose convergence properties are well understood in terms of functional inequalities. In contrast, the analysis of Andrieu et al. (2024) is carried out directly at the level of the discrete-time Markov chain, without relying on an explicit comparison with a continuous-time diffusion. For the Gibbs sampler, this distinction is especially relevant, as there does not appear to exist a natural continuous-time process against which the algorithm can be compared. It is therefore interesting to investigate whether the framework developed by Andrieu et al. (2024) can be applied to the Gibbs sampler targeting non-log-concave distributions, specifically those satisfying either a Poincaré or log-Sobolev inequality.
Another central notion in Andrieu et al. (2024) is that of a so-called close coupling condition for a Markov chain, which provides a quantitative link between geometry and probability. Informally, it requires that whenever two points are “close” in an appropriate metric, the distributions obtained after one step of the Markov chain from these points are “close” in total variation distance. The importance of such connections between geometric and probabilistic distances in the context of mixing times goes back at least to the work of Lovász (1999, Section 5), which studies them in the context of the Hit-and-Run algorithm on convex bodies. In Andrieu et al. (2024), this idea is formalized into a condition that, together with isoperimetry, yields explicit mixing time bounds directly at the level of the discrete-time Markov chain.
@@ 1.1 Our contributions
In this paper, we establish conductance bounds for the systematic-scan and random-scan Gibbs samplers over a broad class of target distributions supported on $\mathbb{R}^{d}$. In particular, consider a target distribution $\pi$ satisfying an $L^{q}$-Poincaré inequality with constant $\mathtt{p}_{q}$ and whose one-dimensional conditional distributions are Hölder continuous in total variation distance with constant $M$ and exponent $\beta$ (see Subsection 2.2 for formal definitions). Denote by $P_{\mathrm{SS}}$ and $P_{\mathrm{RS}}$ the Markov kernels corresponding to the systematic-scan and random-scan Gibbs samplers, respectively. Then, for sufficiently large $d$, our main results (Theorem 4.7 and Corollary 4.9) provide the following bounds on the conductance of the lazy systematic-scan Gibbs kernel, $\Phi(\widetilde{P}_{\mathrm{SS}})$, and the random-scan Gibbs kernel, $\Phi(P_{\mathrm{RS}})$:
$\Phi(\widetilde{P}_{\mathrm{SS}})\gtrsim\frac{\mathtt{p}_{q}}{2^{2q}(Md)^{q/\beta}},$
$\Phi(P_{\mathrm{RS}})\gtrsim\frac{\mathtt{p}_{q}}{2^{2q}(3M)^{q/\beta}(d\cdot\log d)^{q/\beta+1}},$
where $\gtrsim$ denotes inequalities that hold up to a universal constant. Bounds on the mixing time in total variation distance then follow from results of Lovász and Simonovits (1993).
@@
  • We prove new three-set isoperimetric inequalities for distributions satisfying an $L^{q}$-Poincaré or $L^{q}$-log-Sobolev inequality, without assuming log-concavity (Propositions 4.1 and 4.3). This is achieved by applying these functional inequalities to a carefully chosen test function.
@@
  • We show that if $\pi$ is log-Lipschitz continuous or log-smooth, then the one-dimensional conditional distributions under $\pi$ are Hölder continuous in total variation distance (Lemmas 5.4 and 5.5). In the log-smooth case, this is shown under the additional assumption that the one-dimensional conditional distributions of $\pi$ satisfy a uniform $L^{2}$-Poincaré inequality. The proof relies on the fact that log-Lipschitz continuity and log-smoothness (under the above assumption) are inherited by the marginal distributions of $\pi$, as shown in Lemma 5.3.
@@ 1.2 Connections with recent work
As a special case, the conductance approach has been successfully applied to the random-scan Gibbs sampler targeting uniform distributions over convex bodies, where quantitative convergence guarantees have been established (Narayanan and Srivastava, 2022; Laddha and Vempala, 2023; Narayanan et al., 2024). Narayanan and Srivastava (2022) establish a close coupling condition for a multi-step Metropolis-within-Gibbs scheme, where the corresponding Markov kernel updates a randomly chosen coordinate using a single step of random-walk Metropolis while keeping the others fixed. They then compare the conductance of this auxiliary Markov chain with that of the original random-scan Gibbs sampler. Laddha and Vempala (2023) take a different approach, proving a three-set isoperimetric inequality for partitions of space into axis-disjoint regions. This allows them to control the mixing behavior without relying on auxiliary chains.
More recently, Narayanan et al. (2024) refine these results by reducing the dependence on the initialization (the “warm start” parameter) of the mixing time bounds from polynomial to logarithmic. Their approach decomposes the convex body into axis-aligned dyadic cubes and establishes an isoperimetric inequality with respect to a metric that magnifies distances near the boundary. The dimension dependence of their mixing time bounds can be further improved using the sharper isoperimetric inequalities of Fernandez (2025). In this work, we develop a complementary approach for target distributions supported on $\mathbb{R}^{d}$ and establish three-set isoperimetric inequalities with respect to the Euclidean metric that hold for arbitrary partitions. However, it is unclear whether the results in Narayanan and Srivastava (2022); Laddha and Vempala (2023); Narayanan et al. (2024) for sampling from convex bodies can be recovered from our framework.
Additionally, the conductance approach has been used to establish rapid mixing of the Gibbs sampler targeting the posterior distributions in Bayesian probit, logit, and lasso regression models (Lee and Zhang, 2024). In each case, the Gibbs sampler is made tractable by augmenting the state space with latent variables. For such data-augmented chains, Lee and Zhang (2024) show that close coupling conditions follow directly from the conditional independence structure of the posterior distribution. Isoperimetric inequalities are obtained via transference inequalities to a transformed Markov chain with a log-concave target distribution. Under suitable assumptions on the data-generating process, these authors then establish bounds on the mixing time that hold with high probability.
@@ 1.3 Organization
The rest of this paper is structured as follows. Section 2 introduces the Gibbs sampler and outlines the relevant assumptions on the target distribution. Section 3 reviews standard definitions and key results on Markov chain mixing times. In particular, we explain how a three-set isoperimetric inequality and a close coupling condition can be used to derive lower bounds on conductance. Next, we establish three-set isoperimetric inequalities for target distributions satisfying a Poincaré or log-Sobolev inequality (Subsection 4.1) and derive close coupling conditions for the systematic-scan and random-scan Gibbs kernels (Subsection 4.2). We then use these results to prove conductance and mixing time bounds for both variants of the Gibbs sampler (Subsection 4.3). Finally, Section 5 extends these bounds to log-Lipschitz continuous and log-smooth target distributions. We also provide examples of non-log-concave target distributions for which our results can be applied to establish mixing time bounds for both variants of the Gibbs sampler, filling a gap in the existing literature.
@@ 2.1 The Gibbs sampler
We consider a probability distribution $\pi(\mathrm{d}x)$ supported on $\mathbb{R}^{d}$, equipped with its standard Borel $\sigma$-algebra. Let $P(x,\mathrm{d}y)$ be the Markov transition kernel of a discrete-time Markov chain on $\mathbb{R}^{d}$ with invariant distribution $\pi$. Here, the Markov chain is interpreted as the MCMC method for generating approximate samples from $\pi$. The kernel $P$ is said to be reversible with respect to $\pi$ if
@@
Reversibility with respect to $\pi$ implies that $\pi$ is invariant, however the reverse implication is in general false. Among MCMC methods, we focus on the Gibbs sampler. This method has two commonly used variants: the random-scan Gibbs sampler and the systematic-scan Gibbs sampler, which we now introduce. Let $x=\left(x_{1},\dots,x_{d}\right)\in\mathbb{R}^{d}$. For $1\leq k\leq l\leq d$, we write
@@
We additionally use the notation $\left[d\right]=\left\{1,\dots,d\right\}$. Let $X\sim\pi$, that is, $X$ is distributed according to $\pi$. Denote the conditional distribution of $X_{k}$ given $X_{-k}=x_{-k}$ by $\pi(\cdot\mid x_{-k})$. For all $k\in[d]$, define the Markov kernel
$P_{k}(x,\mathrm{d}y)=\pi(\mathrm{d}y_{k}\mid x_{-k})\delta_{x_{-k}}(\mathrm{d}y_{-k}),$ (2)
where $\delta_{x_{-k}}$ denotes the (d1)(d-1)-dimensional Dirac delta distribution: for all measurable $S\subseteq\mathbb{R}^{d-1}$,
$\delta_{x_{-k}}(S)=\begin{cases}1&\mathrm{if}~x_{-k}\in S,\\ 0&\mathrm{if}~x_{-k}\notin S.\end{cases}$
Using the kernels $(P_{k})_{k\in[d]}$, we now define the Markov kernels corresponding to the random-scan and systematic-scan Gibbs samplers. For two Markov kernels $P$ and $Q$, we denote their composition by $PQ(x,\mathrm{d}y)=\int Q(u,\mathrm{d}y)P(x,\mathrm{d}u)$. With this convention, $PQ$ corresponds to first updating the chain according to $P$ and then according to $Q$.
@@
The Markov kernel associated with the random-scan Gibbs sampler is defined as the mixture of the Markov kernels $(P_{k})_{k\in[d]}$:
@@
The Markov kernel associated with the systematic-scan Gibbs sampler is defined as the composition of the Markov kernels $(P_{k})_{k\in[d]}$:
$P_{\mathrm{SS}}=P_{1}P_{2}\dots P_{d-1}P_{d}.$
@@
Naturally, both variants of the Gibbs sampler are $\pi$-invariant. However, the Markov kernel associated with the random-scan Gibbs sampler is reversible with respect to $\pi$, whereas the Markov kernel associated with the systematic-scan Gibbs sampler is generally not. To address this, one approach is to use the multiplicative reversibilization of the systematic-scan Gibbs sampler. The kernel for this version is defined as
$P_{\mathrm{SS}}P_{\mathrm{SS}}^{*}=P_{1}P_{2}\dots P_{d-1}P_{d}^{2}P_{d-1}\dots P_{2}P_{1},$
where $P^{*}$ represents the adjoint of a Markov kernel $P$ (see Subsection 3.1). In this version, the coordinates are updated first in increasing order, from $1$ to $d$, and then in decreasing order, from $d$ to $1$.